OpenStudy (anonymous):

solve please: 29t/8 + t/2 =5

OpenStudy (anonymous):

t = 40/33

OpenStudy (llort):

http://www.wolframalpha.com/input/?i=29t%2F8+%2B+t%2F2+%3D5 Enjoy ;)

OpenStudy (anonymous):

what are the steps?

OpenStudy (anonymous):

first, u got: 29t/8 + t/2 = 5 then, u find a common denominator between the two fractions: 29t/8 + 4t/8 = 5 then, u add the numerators: 33t/8 = 5 then, u multiply both sides by 8: 33t = 5 * 8 = 40 then, u divide both sides by 33, giving u your answer: t = 40/33

OpenStudy (anonymous):

thanks!!